# A Versatility Measure for Parametric Risk Models

### v1

@@ 2.1 The Scientific Method
where $F_{0}\equiv F_{W\mid\boldsymbol{\alpha}}^{\left(0\right)}\left(w\right)$ represents the cumulative distribution function of either $W=X\in\mathbb{\mathbb{Z}}_{0}^{+}$ or $W=Y\in\mathbb{R}_{0}^{+}$, with unknown $k$-dimensional parameter vector $\boldsymbol{a}=\left[a_{1},\ldots,a_{k}\right]$. The following subsections address characteristics of $F_{0}$ that are particularly desirable in this setting.
@@ 2.2 Parsimony and Functional Simplicity
To clarify these relationships, let the real-valued quantities $\mathcal{I}\left(F_{0}\right)\geq 0$, $\mathcal{C}\left(F_{0}\right)\geq 0,$ and $\mathcal{T}\left(F_{0}\right)\geq 0$ denote hypothetical heuristic measures of the intrinsic idiosyncrasy, complexity, and contrivance of the distribution function $F_{0}$, where each quantity is assumed to be independent of the number of parameters, $k$. In addition, set $\mathcal{V}\left(F_{0}\right)=1/\mathcal{I}\left(F_{0}\right),$ $\mathcal{S}\left(F_{0}\right)=1/\mathcal{C}\left(F_{0}\right),$ and $\mathcal{A}\left(F_{0}\right)=1/\mathcal{T}\left(F_{0}\right)$ as the corresponding inverse (i.e., opposite) quantities: versatility, simplicity, and adaptability, respectively. Then, from the discussion of the preceding paragraph, it follows (heuristically) that
$\mathcal{I}\left(F_{0}\right)=\mathcal{I}_{1}\left(\mathcal{C}^{\left(+\right)},\mathcal{T}^{\left(+\right)}\right)$
$=\mathcal{I}_{2}\left(\mathcal{S}^{\left(-\right)},\mathcal{A}^{\left(-\right)}\right),$ (2)
where the superscripts indicate the direction of the argument’s relationship with $\mathcal{I}$ (i.e., $\left(+\right)=\textrm{positive}$ or $\left(-\right)=\textrm{negative}$).
Now let $\mathcal{P}\left(F_{0}\right)\geq 0$ denote a hypothetical real-valued measure of the parsimony of $F_{0}$, taken to be negatively related to both its number of parameters and functional complexity (because both quantities oppose the objective of simplicity), but independent of functional contrivance (because unnaturally manipulated functions can possess simple or complex structures). We then can write (heuristically)
$\mathcal{P}\left(F_{0}\right)=\mathcal{P}_{1}\left(k^{\left(-\right)},\mathcal{C}^{\left(-\right)}\right)$
$=\mathcal{P}_{2}\left(k^{\left(-\right)},\mathcal{S}^{\left(+\right)}\right).$ (3)
@@ 2.3 Robustness and Functional Adaptability
Letting $\mathcal{R}\left(F_{0}\right)\geq 0$ denote a hypothetical real-valued measure of the robustness of $F_{0}$, we will take this quantity to be positively related to its number of parameters (more of which permit greater accuracy of fit), negatively related to its degree of functional contrivance (because unnaturally manipulated functions are less adaptable to a wide range of empirical observations), and independent of functional complexity (because unnaturally complicated functions can enjoy various degrees of adaptability). We thus write (again heuristically)
$\mathcal{R}\left(F_{0}\right)=\mathcal{R}_{1}\left(k^{\left(+\right)},\mathcal{T}^{\left(-\right)}\right)$
$=\mathcal{R}_{2}\left(k^{\left(+\right)},\mathcal{A}^{\left(+\right)}\right).$ (4)
@@ 2.4 Functional Versatility
With this in mind, we treat functional versatility as the inverse of idiosyncrasy:
$\mathcal{V}\left(F_{0}\right)=\mathcal{V}\left(\mathcal{S}^{\left(+\right)},\mathcal{A}^{\left(+\right)}\right),$ (5)
@@
Simplicity $\mathcal{S}\left(F_{0}\right)=1/\mathcal{C}\left(F_{0}\right)$ Complexity
Parsimony $\mathcal{P}\left(F_{0}\right)=\mathcal{P}_{2}\left(\mathcal{S}^{\left(+\right)},k^{\left(-\right)}\right)$
$=\mathcal{P}_{1}\left(\mathcal{C}^{\left(-\right)},k^{\left(-\right)}\right)$
Adaptability $\mathcal{A}\left(F_{0}\right)=1/\mathcal{T}\left(F_{0}\right)$ Contrivance
Robustness $\mathcal{R}\left(F_{0}\right)=\mathcal{R}_{2}\left(\mathcal{A}^{\left(+\right)},k^{\left(+\right)}\right)$
$=\mathcal{R}_{1}\left(\mathcal{T}^{\left(-\right)},k^{\left(+\right)}\right)$
@@
$=1/\mathcal{I}\left(F_{0}\right)$
@@ 3.1 Two Types of Complexity
With regard to Type A, it is important to note that the complexity of a PMF/PDF’s notational form should not be conflated with notions of artificial (and possibly tendentious) design, even if we know or suspect a particular PMF/PDF to have been constructed by a clever probabilist with specific objectives in mind. Certainly, functions can be devised to achieve high levels of Kolmogorov (algorithmic) complexity (also called Kolmogorov entropy), meaning that they cannot be compressed into short string lengths in a mathematical coding system. [^8] See Li and Vitányi (2019) for a recent treatment of Kolmogorov complexity/entropy. However, such constructions necessitate intentionally injecting elements of disorder or disorganization, since functions formed according to an ordered plan or pattern (even if somewhat complicated) possess less Kolmogorov complexity because they are more compressible.
In the case of Type B, we encounter an apparent paradox: PMFs/PDFs with greater Shannon (information-theoretic) entropy (also called Shannon information) tend to possess simpler, flatter shapes. [^9] See Gray (2011) for a comprehensive overview of Shannon entropy. For example, the maximum-entropy discrete distribution on $x\in\mathbb{\mathbb{Z}}_{0}^{+}$, subject to a fixed mean, is given by the $\textrm{Geometric}\left(\varrho\right)$ PMF ($f_{X\mid\varrho}\left(x\right)=\varrho\left(1-\varrho\right)^{x}$, for $\varrho\in\left(0,1\right)$), and the maximum-entropy continuous distribution on $y\in\mathbb{R}_{0}^{+}$, again subject to a fixed mean, is given by the $\textrm{Exponential}\left(\lambda\right)$ PDF ($f_{Y\mid\lambda}\left(y\right)=\lambda e^{-\lambda y}$, for $\lambda\in\mathbb{R}^{+}$). This is attributable to the fact that, for a given sample space and specified constraints (e.g., a fixed mean or other moment), PMFs/PDFs providing greater Shannon entropy, as measured by the expected “surprise” from a random trial, are those that spread the distribution’s probability more evenly across the sample space. Essentially, we can say (informally, at least) that increasing the disorder or disorganization of random draws is achieved by decreasing the complexity of the distribution function’s shape.
@@ 4.1 Fisher Information
For the moment, assume $F_{0}\equiv F_{W\mid a}\left(w\right)$, where $a\in\mathbb{R}^{+}$ is a scalar parameter. To evaluate the functional adaptability of this probability distribution, it is natural to consider the impact of $a$ on the corresponding PMF or PDF, $f_{0}\left(w\right)$ – an effect that is measured by the Fisher information associated with a single observation of $W$:
$\textrm{I}_{F_{0}}\left(a\right)=\textrm{E}_{F_{0}}\left[\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a}\right)^{2}\right]$ (11)
$=-\textrm{E}_{F_{0}}\left[\dfrac{\partial^{2}\ln\left(f_{0}\left(W\right)\right)}{\partial a^{2}}\right].$
At a given value of $a$, the above equations quantify the extent by which changes in the parameter modify the shape of the PMF or PDF, on average. [^12] See Lehmann and Casella (1998) for a detailed exposition of Fisher information.
Like Shannon entropy, Fisher information is additive for independent random variables. [^13] Although Fisher information and Shannon entropy (information) possess certain similarities, they are designed to measure different types of “information” in different contexts, and therefore are not directly comparable. Nevertheless, they do share a mathematical connection through the calculation of relative entropy (Kullback-Leibler divergence). See, e.g., Cover and Thomas (2006). In particular, for a sample of $n$ independent and identically distributed (IID) observations, $\boldsymbol{W}=\left[W_{1},\ldots,W_{n}\right]$, (11) aggregates to
@@
[…]
@@
to become much more sensitive to errors in the parameter estimator. For example, assuming an error of $+0.10$ in $\hat{\lambda}_{n}$, we would estimate the probability that $Y$ exceeds 1 as $0.0022$, which is fairly close to the true value of approximately $0.0025$. However, assuming the same error of $+0.10$ in $\hat{\theta}_{n}$, yields a corresponding probability estimate of $0.0235$, which is substantially farther from the true value.
@@ 4.2 $\boldsymbol{k^{\textrm{th}}}$ Root of Unnormalized Jeffreys Prior
$\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)=\textrm{E}_{F_{0}}\left[\begin{array}[]{ccc}\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)^{2}&\cdots&\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)\\ \vdots&\ddots&\vdots\\ \left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)&\cdots&\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)^{2}\end{array}\right].$ (14)
Since the determinant of this matrix represents the volume spanned by the parallelepiped formed by the $k$ columns (or rows), it constitutes a natural extension to $k$ dimensions of the scalar measure of the average impact of $a$ on the PMF/PDF in (11). One then can raise this quantity to the $\tfrac{1}{2k}$ power to: (i) reduce its dimensional units from $k$ to 1, allowing it to be expressed on a “per-parameter” basis; and (ii) return the squared units formed by all matrix elements in (14) to their original scale (in the same way exponentiating the Shannon entropy did in Section 3). Interestingly, the resulting quantity, $\left(\det\left(\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)\right)\right)^{1/2k}$, is equivalent to the $k^{\textrm{th}}$ root of the unnormalized Jeffreys prior PMF/PDF for the parameter vector $\boldsymbol{a}=\left[a_{1},\ldots,a_{k}\right]$:

### v3

@@ 2.1 The Scientific Method
where $F_{0}\equiv F_{W\mid\boldsymbol{a}}^{\left(0\right)}\left(w\right)$ represents the cumulative distribution function of either $W=X\in\mathbb{\mathbb{Z}}_{0}^{+}$ or $W=Y\in\mathbb{R}_{0}^{+}$, with unknown $k$-dimensional parameter vector $\boldsymbol{a}=\left[a_{1},\ldots,a_{k}\right]$. The following subsections address characteristics of $F_{0}$ that are particularly desirable in this setting.
@@ 2.2 Parsimony and Functional Simplicity
To clarify these relationships, let the real-valued quantities $\mathcal{I}\left(F_{0}\right)\geq 0$, $\mathcal{C}\left(F_{0}\right)\geq 0,$ and $\mathcal{T}\left(F_{0}\right)\geq 0$ denote hypothetical heuristic measures of the intrinsic idiosyncrasy, complexity, and contrivance of the distribution function $F_{0}$, where each quantity is assumed to be independent of the number of parameters, $k$. In addition, set $\mathcal{V}\left(F_{0}\right)=\tfrac{1}{\mathcal{I}\left(F_{0}\right)},$ $\mathcal{S}\left(F_{0}\right)=\tfrac{1}{\mathcal{C}\left(F_{0}\right)},$ and $\mathcal{A}\left(F_{0}\right)=\tfrac{1}{\mathcal{T}\left(F_{0}\right)}$ as the corresponding inverse (i.e., opposite) quantities: versatility, simplicity, and adaptability, respectively. Then, from the discussion of the preceding paragraph, it follows that $\mathcal{I}\left(F_{0}\right)$ may be viewed (heuristically) as a bivariate function of $\mathcal{C}\left(F_{0}\right)$ and $\mathcal{T}\left(F_{0}\right)$ (or alternatively, of $\mathcal{S}\left(F_{0}\right)$ and $\mathcal{A}\left(F_{0}\right)$); that is,
$\mathcal{I}\left(F_{0}\right)=\mathcal{I}\left(\mathcal{C}^{\left(+\right)},\mathcal{T}^{\left(+\right)}\right)$
$=\mathcal{I}\left(\mathcal{S}^{\left(-\right)},\mathcal{A}^{\left(-\right)}\right),$ (2)
where the superscripts indicate the direction of the argument’s impact on $\mathcal{I}$ (i.e., $\left(+\right)=\textrm{positive}$ and $\left(-\right)=\textrm{negative}$).
Now let $\mathcal{P}\left(F_{0}\right)\geq 0$ denote a hypothetical real-valued measure of the parsimony of $F_{0}$, taken to be negatively related to both its number of parameters and functional complexity (because both quantities oppose the objective of simplicity), but independent of functional contrivance (because unnaturally manipulated functions can possess simple or complex structures). We then can treat $\mathcal{P}\left(F_{0}\right)$ (heuristically) as a bivariate function of $k$ and $\mathcal{C}\left(F_{0}\right)$ (or alternatively, of $k$ and $\mathcal{S}\left(F_{0}\right)$):
$\mathcal{P}\left(F_{0}\right)=\mathcal{P}\left(k^{\left(-\right)},\mathcal{C}^{\left(-\right)}\right)$
$=\mathcal{P}\left(k^{\left(-\right)},\mathcal{S}^{\left(+\right)}\right){\color[rgb]{1,0,0},}$ (3)
where the superscripts again indicate the direction of the argument’s impact.
@@ 2.3 Robustness and Functional Adaptability
Letting $\mathcal{R}\left(F_{0}\right)\geq 0$ denote a hypothetical real-valued measure of the robustness of $F_{0}$, we will take this quantity to be positively related to its number of parameters (more of which permit greater accuracy of fit), negatively related to its degree of functional contrivance (because unnaturally manipulated functions are less adaptable to a wide range of empirical observations), and independent of functional complexity (because unnaturally complicated functions can enjoy various degrees of adaptability). We thus can view $\mathcal{R}\left(F_{0}\right)$ (heuristically) as a bivariate function of $k$ and $\mathcal{T}\left(F_{0}\right)$ (or alternatively, of $k$ and $\mathcal{A}\left(F_{0}\right)$):
$\mathcal{R}\left(F_{0}\right)=\mathcal{R}\left(k^{\left(+\right)},\mathcal{T}^{\left(-\right)}\right)$
$=\mathcal{R}\left(k^{\left(+\right)},\mathcal{A}^{\left(+\right)}\right).$ (4)
@@ 2.4 Functional Versatility
With this in mind, we will focus on functional versatility as the inverse of idiosyncrasy:
$\mathcal{V}\left(F_{0}\right)=\dfrac{1}{\mathcal{I}\left(\mathcal{S}^{\left(-\right)},\mathcal{A}^{\left(-\right)}\right)}$
$=\mathcal{V}\left(\mathcal{S}^{\left(+\right)},\mathcal{A}^{\left(+\right)}\right),$ (5)
@@
Simplicity $\mathcal{S}\left(F_{0}\right)=\dfrac{1}{\mathcal{C}\left(F_{0}\right)}$ Complexity
Parsimony $\mathcal{P}\left(F_{0}\right)=\mathcal{P}\left(k^{\left(-\right)},\mathcal{S}^{\left(+\right)}\right)$
$=\mathcal{P}\left(k^{\left(-\right)},\mathcal{C}^{\left(-\right)}\right)$
Adaptability $\mathcal{A}\left(F_{0}\right)=\dfrac{1}{\mathcal{T}\left(F_{0}\right)}$ Contrivance
Robustness $\mathcal{R}\left(F_{0}\right)=\mathcal{R}\left(k^{\left(+\right)},\mathcal{A}^{\left(+\right)}\right)$
$=\mathcal{R}\left(k^{\left(+\right)},\mathcal{T}^{\left(-\right)}\right)$
@@
$=\dfrac{1}{\mathcal{I}\left(F_{0}\right)}$
@@ 3.1 Two Types of Complexity
With regard to Type A, it is important to note that the complexity of a PMF/PDF’s notational form should not be conflated with notions of artificial (and possibly tendentious) design, even if we know or suspect a particular PMF/PDF to have been constructed by a clever probabilist with specific objectives in mind. Certainly, functions can be devised to achieve high levels of Kolmogorov (algorithmic) complexity (also called Kolmogorov entropy), meaning that they cannot be compressed into short string lengths in a mathematical coding system. [^8] See Li and Vitányi (2019) for a general overview of Kolmogorov complexity/entropy and Lu and Oliveira (2022) for insights into current applications. However, such constructions necessitate intentionally injecting elements of disorder or disorganization, since functions formed according to an ordered plan or pattern (even if somewhat complicated) possess less Kolmogorov complexity because they are more compressible.
In the case of Type B, we encounter an apparent paradox: PMFs/PDFs with greater Shannon (information-theoretic) entropy (also called Shannon information) tend to possess simpler, flatter shapes. [^9] See Gray (2011) for a broad exposition of Shannon entropy and Nanda and Chowdhury (2021) for a survey of major developments up to the present. For example, the maximum-entropy discrete distribution on $x\in\mathbb{\mathbb{Z}}_{0}^{+}$, subject to a fixed mean, is given by the $\textrm{Geometric}\left(\varrho\right)$ PMF ($f_{X\mid\varrho}\left(x\right)=\varrho\left(1-\varrho\right)^{x}$, for $\varrho\in\left(0,1\right)$), and the maximum-entropy continuous distribution on $y\in\mathbb{R}_{0}^{+}$, again subject to a fixed mean, is given by the $\textrm{Exponential}\left(\lambda\right)$ PDF ($f_{Y\mid\lambda}\left(y\right)=\lambda e^{-\lambda y}$, for $\lambda\in\mathbb{R}^{+}$). This is attributable to the fact that, for a given sample space and specified constraints (e.g., a fixed mean or other moment), PMFs/PDFs providing greater Shannon entropy, as measured by the expected “surprise” from a random trial, are those that spread the distribution’s probability more evenly across the sample space. Essentially, we can say (informally, at least) that increasing the disorder or disorganization of random draws is achieved by decreasing the complexity of the distribution function’s shape.
@@ 4.1 Fisher Information
For the moment, assume $F_{0}\equiv F_{W\mid a}^{\left(0\right)}\left(w\right)$, where $a$ is a scalar parameter belonging to some open set $\mathbb{A}\subseteq\mathbb{R}$. To evaluate the functional adaptability of this probability distribution, it is natural to consider the impact of $a$ on the corresponding PMF or PDF, $f_{0}\left(w\right)$ – an effect that is measured by the Fisher information associated with a single observation of $W$:
$\textrm{I}_{F_{0}}\left(a\right)=\textrm{E}_{F_{0}}\left[\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a}\right)^{2}\right]$
$=-\textrm{E}_{F_{0}}\left[\dfrac{\partial^{2}\ln\left(f_{0}\left(W\right)\right)}{\partial a^{2}}\right],$ (11)
[…]
Like Shannon entropy, Fisher information is additive for independent random variables. [^13] Although Fisher information and Shannon entropy (information) possess certain similarities, they are designed to measure different types of “information” in different contexts, and therefore are not directly comparable. Nevertheless, they do share a mathematical connection through the calculation of relative entropy (Kullback-Leibler divergence). See, e.g., Cover and Thomas (2006). In particular, for a sample of $n$ independent and identically distributed (IID) observations, $\boldsymbol{W}=\left[W_{1},W_{2},\ldots,W_{n}\right]$, (11) aggregates to
@@
The interplay of negative and positive implications of the Fisher information becomes apparent in the context of the Cramér-Rao lower bound on the asymptotic variance of any consistent estimator, $\hat{a}_{n}$, of $a$: [^14] A consistent estimator, $\hat{a}_{n}$, converges in probability to the true parameter, $a$. This is denoted by $\hat{a}_{n}\overset{\mathcal{P}}{\rightarrow}a$; that is, for any $\epsilon>0$, $\underset{n\rightarrow\infty}{\lim}\Pr\left\{\left|\hat{a}_{n}-a\right|>\epsilon\right\}=0$.
@@
to become much more sensitive to errors in the parameter estimator. For example, assuming an error of $+0.1000$ in $\hat{\lambda}_{n}$, we would estimate $\Pr\left\{Y>1\right\}$ as $0.0022$, which falls short of the true value ($\Pr\left\{Y>1\right\}\approx 0.0025$) by $0.0003$. To achieve a comparably small deviation in the estimate of $\Pr\left\{Y>1\right\}$ under the alternative parameterization would require the absolute error in $\hat{\theta}_{n}$ be no greater than
$\max\left\{\left|\dfrac{1}{-\ln\left(0.0025-0.0003\right)}-\dfrac{1}{6}\right|,\left|\dfrac{1}{-\ln\left(0.0025+0.0003\right)}-\dfrac{1}{6}\right|\right\}\approx\max\left\{\left|-0.0032\right|,\left|0.0035\right|\right\}$
$=0.0035.$
Computation of the Fisher information frequently arises in the context of maximum-likelihood estimation (MLE) because the estimator generated, $\hat{a}_{n}$, is (under the regularity conditions identified in Footnote 12) asymptotically Normal, [^15] An asymptotically Normal estimator, $\hat{a}_{n}$, converges in distribution to the Normal distribution with mean $a$ (the true parameter) and variance equal to the Cramér-Rao lower bound. This is denoted by $\hat{a}_{n}\overset{\mathcal{D}}{\rightarrow}\textrm{Normal}\left(a,\>\tfrac{1}{\textrm{I}_{F_{0}}\left(a,n\right)}\right)$; that is, $\underset{n\rightarrow\infty}{\lim}\left|\Pr\left\{\tfrac{\hat{a}_{n}-a}{\sqrt{1/\textrm{I}_{F_{0}}\left(a,n\right)}}\leq t\right\}-\Phi\left(t\right)\right|=0$ for all $t\in\mathbb{R}$, where $\Phi\left(\cdot\right)$ is the CDF of the Standard Normal distribution. permitting its variance to be estimated well by the inverse of the Fisher information; that is,
$\textrm{Var}{}_{F_{0}}\left[\hat{a}_{n}\right]\approx\dfrac{1}{\textrm{I}_{F_{0}}\left(\hat{a}_{n},n\right)}$
for large $n,$ where $\textrm{I}_{F_{0}}\left(\hat{a}_{n},n\right)$ denotes an estimator of $\textrm{I}_{F_{0}}\left(a,n\right)$ obtained by substituting $\hat{a}_{n}$ for $a$. Two common obstacles to employing the Fisher information in this way are: (1) the failure of $W\sim F_{0}\left(w\right)$ to satisfy the first condition in Footnote 12 because one or both endpoints of the sample space ($\mathbb{W}$) depend on $a$; and (2) the non-concavity of the log-likelihood, ${\textstyle\sum_{i=1}^{n}}\ln\left(f_{0}\left(W_{i}\right)\right)$, as a function of $a$.
@@ 4.2 $\boldsymbol{k^{\textrm{th}}}$ Root of Unnormalized Jeffreys Prior
$\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)=\textrm{E}_{F_{0}}\left[\begin{array}[]{ccc}\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)^{2}&\cdots&\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)\\ \vdots&\ddots&\vdots\\ \left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{1}}\right)&\cdots&\left(\dfrac{\partial\ln\left(f_{0}\left(W\right)\right)}{\partial a_{k}}\right)^{2}\end{array}\right]{\color[rgb]{1,0,0},}$ (14)
which is well defined subject to natural extensions of the regularity conditions stated in Footnote 12.
Since the determinant of this matrix represents the volume spanned by the parallelepiped formed by the $k$ columns (or rows), it constitutes a natural extension to $k$ dimensions of the scalar measure of the average impact of $a$ on the PMF/PDF in (11). One then can raise this quantity to the $\tfrac{1}{2k}$ power to: (i) reduce its dimensional units from $k$ to 1, allowing it to be expressed on a “per-parameter” basis; and (ii) return the squared units formed by all matrix elements in (14) to their original scale (in the same way exponentiating the Shannon entropy did in Section 3). Interestingly, the resulting quantity, $\left(\det\left(\mathbf{I}_{F_{0}}^{\left(k\right)}\left(\boldsymbol{a}\right)\right)\right)^{1/2k}$, is equivalent to the $k^{\textrm{th}}$ root of the unnormalized Jeffreys prior PMF/PDF for the parameter vector $\boldsymbol{a}=\left[a_{1},a_{2},\ldots,a_{k}\right]$: